# GraphPro: Graph Pre-training and Prompt Learning for Recommendation

### 20 Dec 2023 v2

@@ Abstract.
GNN-based recommenders have excelled in modeling intricate user-item interactions through multi-hop message passing. However, existing methods often overlook the dynamic nature of evolving user-item interactions, which impedes the adaption to changing user preferences and distribution shifts in newly arriving data. Thus, their scalability and performances in real-world dynamic environments are limited. In this study, we propose GraphPro, a framework that incorporates parameter-efficient and dynamic graph pre-training with prompt learning. This novel combination empowers GNNs to effectively capture both long-term user preferences and short-term behavior dynamics, enabling the delivery of accurate and timely recommendations. Our GraphPro framework addresses the challenge of evolving user preferences by seamlessly integrating a temporal prompt mechanism and a graph-structural prompt learning mechanism into the pre-trained GNN model. The temporal prompt mechanism encodes time information on user-item interaction, allowing the model to naturally capture temporal context, while the graph-structural prompt learning mechanism enables the transfer of pre-trained knowledge to adapt to behavior dynamics without the need for continuous incremental training. We further bring in a dynamic evaluation setting for recommendation to mimic real-world dynamic scenarios and bridge the offline-online gap to a better level. Our extensive experiments including a large-scale industrial deployment showcases the lightweight plug-in scalability of our GraphPro when integrated with various state-of-the-art recommenders, emphasizing the advantages of GraphPro in terms of effectiveness, robustness and efficiency.
copyright: none
@@ 1. Introduction
Recommender systems are integral to numerous Web platforms, assisting users in navigating through the overwhelming amount of information by suggesting relevant items (Geng et al., 2022; Wen et al., 2023). In recent years, graph neural networks (GNNs) have emerged as powerful tools for modeling user-item interactions in recommendation tasks, enabling effective representation learning on graph-structured data. By treating users and items as nodes and their interactions as edges, GNNs can capture intricate multi-hop relationships between users and items, facilitating the generation of personalized recommendations.
Earlier works (Berg et al., 2017; Ying et al., 2018; Wang et al., 2019) on GNN-enhanced recommendation have primarily focused on designing effective message passing mechanisms to capture collaborative relations between users and items. These works aim to leverage the power of GCN to capture high-order connectivity in the user-item interaction graph. Subsequent research has further explored simplifying the message passing process (Chen et al., 2020a; He et al., 2020), reducing the complexity of GNN models (Mao et al., 2021; Shen et al., 2021), and improving the quality of sampling techniques (Huang et al., 2021). More recently, researchers have advanced graph-based recommenders by incorporating self-supervised learning (SSL) techniques into GNNs (Yu et al., 2023). These methods (Wu et al., 2021; Cai et al., 2023a; Yu et al., 2022) generally employ the InfoNCE (Oord et al., 2018) loss to align contrastive views, thereby denoising and improving the robustness of the base LightGCN (He et al., 2020) model.
While these methods have shown impressive performance, they have primarily focused on static scenarios (Figure 1 upper), overlooking the dynamic nature of recommendation, where new user-item interactions continue to evolve over time, often with distribution shifts reflecting changing user interests (Xiao et al., 2023; Yang et al., 2023). In real-world scenarios (Figure 1 lower), it exhibits a dynamic setting, where the model recursively learns from newly arriving data, and predicts for current time. However, existing methods primarily design a model for single-graph training and evaluation, which leads to a degradation in recommendation dynamics and widens the offline-online gap (Ji et al., 2023). Additionally, the arrival of new data may exhibit distribution shifts, further complicating the task of making accurate recommendations for graph-based recommendation models without incorporating useful contextual information for the newly arrived data. These challenges significantly limit the scalability of existing models and hinder their ability to adapt to evolving user preferences in a timely manner. This is crucial for providing up-to-date and accurate recommendations in dynamic environments.
@@
To tackle the challenges, we propose a simple yet effective framework named GraphPro, which integrates parameter-efficient and dynamic Graph Pre-Training with Prompt Learning for recommendation. Our proposed method entails pre-training GNNs on extensive historical interaction data, followed by fine-tuning them on more recent target data using time-aware graph prompt learning. In the pre-training phase, the model assimilates knowledge from a substantial amount of historical interactions, effectively capturing long-term user preferences and item relevance. Subsequently, during the fine-tuning phase on the target time periods of data, the model swiftly adapts to evolving user preferences and captures short-term behavior dynamics. This is achieved through a prompt learning schema, facilitating effective knowledge transfer.
@@
To ensure that the pre-trained GNN effectively handles evolving user preferences, our GraphPro framework seamlessly integrates a temporal prompt mechanism alongside a graph-structural prompt learning mechanism. This integration enables the injection of time-aware context from new data, empowering the model to adapt to changing user preferences. Drawing inspiration from advances in relative positional encoding techniques (Press et al., 2021; Su et al., 2021), we meticulously design a dedicated temporal prompt mechanism that aligns with the message aggregation layer of GNNs. Within this prompt mechanism, we encode time information on interaction edges as part of the normalization term for aggregation, all in a parameter-free manner. This innate capability allows the model to naturally incorporate temporal information without the need for additional fine-tuning. By imbuing the pre-trained graph model with temporal awareness, we empower the model to effectively capture vital signals that are more pertinent to the evolving user preferences.
@@
Furthermore, our graph-structural prompt learning mechanism facilitates the seamless transfer of knowledge from the pre-trained model to downstream short-term and recent recommendation tasks. This framework eliminates the requirement for continuous incremental learning of the pre-trained model, instead enabling the transfer of pre-trained knowledge to any future time period to effectively adapt to behavior dynamics. In this mechanism, we incorporate the newly generated interaction edges between the fine-tuning time and the pre-training time as prompt edges. By including these prompt edges, we provide the pretrained model with essential contextual information for the fine-tuning process. Rather than undergoing extensive training, we perform a single non-training forward pass on the prompt edges. This prompts the pretrained model to adapt to the distribution shift of node representations and effectively adjusts its predictions accordingly. It is worthwhile mentioning that our GraphPro is model-agnostic and parameter-efficient, allowing for seamless integration into existing GNN recommenders as a plug-in enhancement. In summary, the main contributions of our work can be summarized as follows:
@@
  • We emphasize the criticality of effectively and scalably pre-training and fine-tuning graph-based recommenders for time-evolving user preferences, thus facilitating up-to-date and accurate recommendations in dynamic environments.
@@
  • We present GraphPro that effectively handles evolving user preferences by pre-training and fine-tuning GNNs. The proposed prompt learning paradigm facilitates the transfer of valuable and relevant knowledge from the pre-trained model to downstream recommendation tasks in both temporal and structural manners.
@@
  • We further introduce a snapshot-based dynamic setting for recommendation evaluation. Compared to the vanilla single-time testing, it brings better approximation to real-world scenarios.
@@
  • We conduct extensive experiments on diverse datasets to demonstrate the advantages of GraphPro in terms of robustness, efficiency and performance. To further justify the effectiveness of our framework, we include an online industry deployment with A/B testing on a large-scale online platform.
To ensure result reproducibility, we provide the implementation details and source code of our proposed framework at: https://github.com/HKUDS/GraphPro/
@@ 3.1.2. Temporal Prompt-enhanced Graph Convolutions
Here, $t^{\text{unix}}_{u,v}$ denotes the Unix timestamp assigned to the edge $e_{u,v}$, and the $\lfloor$*$\rfloor$ * ⌋ notation denotes the floor operation. To avoid the influence of specific numerical scales and ensure uniformity, we normalize these time attributes $\boldsymbol{t}={t_{u,v}|e_{u,v}=1}$ to the range of [0, 1].
@@
3.2. Fine-Tuning with Graph-Structural
@@ 4. Evaluation
In this section, we compare our proposed GraphPro with state-of-the-art methods across diverse research lines and settings, with the aim of addressing the research questions shown below.
  • RQ1: Can GraphPro outperform state-of-the-art time-aware graph learning models and pre-trained GNNs in making dynamic recommendations across different time slots?
@@
  • RQ3: Can GraphPro perform on par or even outperform the vanilla full-data training paradigm?
@@ 4.1.1. Datasets
We adopt three public datasets covering diverse real-world scenarios of dynamic recommendation. Taobao records implicit feedback from Taobao.com, a Chinese e-commerce platform during 10 days. Koubei dataset, provided for the IJCAI’16 contest, records 9 weeks of user interactions with nearby stores on Koubei in Alipay. Amazon dataset consists of a 13-week’s collection of product reviews sourced from Amazon. The details can be referred in Table 1.
@@ 4.1.2. Baseline Models
We include the recent dynamic graph neural networks and graph prompt approaches as our baselines. Specifically, three most relevant research lines are included for comparison:
Dynamic Recommendation Methods. We include DGCN (Li et al., 2020) which also studies dynamic learning in collaborative filtering. It categorizes edges as past and current ones, and designs a new GNN framework that makes information flow from past edges to current. However, it does not include an explicit dynamic setting with snapshots, and focuses learning on a single graph.
@@
Dynamic Graph Neural Networks. They focus on addressing dynamic graphs by updating previously learned embeddings in a time-aware manner, and handling graph dynamics. As competitors for comparison, we include EvolveGCN (Pareja et al., 2020) and ROLAND (You et al., 2022).
  • EvolveGCN (Pareja et al., 2020). This method addresses the dynamism of graph sequences by utilizing an RNN to adapt the parameters of the Graph Convolutional Network (GCN) over time. The GCN parameters can be either hidden states (referred to as the -H variant) or inputs of a recurrent architecture (referred to as the -O variant).
  • ROLAND (You et al., 2022). This state-of-the-art dynamic graph learning baseline utilizes a meta-learning approach to update previously learned embeddings for re-initialization. These updated embeddings are then fused with layer-wise hidden states of GNN.
@@ 4.1.3. Integration with GNN Recommenders
To highlight its versatility, GraphPro serves as a general architecture that can be seamlessly integrated as a plug-in component with any GNN-based recommender. In our evaluation, we implement GraphPro using the LightGCN (He et al., 2020) model, renowned for its simplicity and efficiency. Furthermore, we extend the applicability of GraphPro by integrating it with SOTA recommenders that incorporate self-supervised learning designs, such as SGL (Wu et al., 2021), MixGCF (Huang et al., 2021), and SimGCL (Yu et al., 2022), providing empirical evidence of the GraphPro’s effectiveness in enhancing them for dynamic and adaptable recommendations.
@@ 4.1.4. Evaluation Protocols
Our evaluation settings encompass graph snapshots of multiple time intervals (e.g., day, week) to further simulate the practical challenges of learning dynamics. To learn from snapshots, we fine-tune all the models using a 2-size sliding window approach to continuously learn from the current snapshot and make predictions for the next. For the proposed P-L paradigm, we take a large portion of the data to pre-train, and use the remaining snapshots for fine-tuning and testing as shown in Table 1. To ensure fair comparison free of data inequality, the baselines also follow the same proposed paradigm. For example, for dynamic GNNs, the very initial weights for tuning are also trained over the pre-training time span. We report metrics averaged over all target future temporal snapshots as following the common setting in (Pareja et al., 2020; You et al., 2022). We report Recall@k and nDCG@k at k=20, following the common all-rank settings in previous works (He et al., 2020; Wu et al., 2021; He et al., 2017).
@@ 4.2.1. Comparison with Baselines
We present the performance of our GraphPro method as well as other alternative solutions, including graph prompt methods and dynamic GNNs, with LightGCN as the base model, as summarized in Table 2. Analyzing the results, we have made the following key observations:
  • Our GraphPro consistently outperforms graph prompt and dynamic graph learning methods, demonstrating the superiority of our pre-training and prompt learning design. Specifically, we observe average improvements of 6.0%, 3.2%, and 8.3% in Recall and nDCG across the three datasets. We attribute these advantages to two key factors: 1) Our temporal prompt mechanism empowers GraphPro to capture dynamically evolving user-item interactions throughout both the pre-training and fine-tuning stages. 2) The graph-structural prompt design facilitates effective knowledge transfer from the large-scale pre-trained model, addressing distribution shift issues between temporal snapshots.
  • The absence of a consistent winner among the baseline methods indicates the challenging nature of the dynamic recommendation task. Despite EvolveGCN’s impressive performance on the Taobao dataset, it does not exhibit a distinct advantage over other models on the remaining datasets. This discrepancy may be due to potential overfitting issues related to short-term user patterns, resulting from the use of complex neural architectures to encode user-item interactions in each new time period. In contrast, our GraphPro utilizes lightweight prompt mechanisms that effectively capture both long-term user interests and incorporate new preference context from recently observed behavior data.
@@ 4.2.2. Integration with SOTA Methods
Additionally, we assess the adaptability of GraphPro across different backbone recommenders, namely MixGCF, SGL, and SimGCL. We re-implement all methods on these base models using the same evaluation settings. The evaluation results, averaged across multiple time slots, are presented in Table 2. Our observations are summarized as follows:
  • GraphPro continues to demonstrate superior performance when integrated with state-of-the-art recommenders. This highlights the remarkable adaptability of our approach, enhancing the performance of different state-of-the-art models in diverse scenarios. Baseline methods exhibit diverse performance rankings across different base recommenders and datasets. No single method consistently outperforms others across all three datasets. On Taobao, EvolveGCN-O is generally on par as the second-best method. However, on Amazon, GPF consistently outperforms other baselines with all three base recommenders. This sensitivity to data characteristics and base models hinders the baseline methods from yielding stable and significant results.
  • In general, better representation learning capabilities in the base model lead to higher performance. When comparing GraphPro’s performance in the SGL+Taobao and SimGCL+Taobao settings, there is a 4.5% improvement. This indicates that our approach effectively benefits from the enhanced representation provided by the base model and performs exceptionally well in dynamic scenarios. However, methods like EvolveGCN-O do not show improved performance as the base model’s representation capabilities increase. This suggests that their designs may lack strong generalization ability and could even yield negative outcomes.

### 14 Jan 2024 v3

@@ Abstract.
GNN-based recommendation systems have excelled at capturing complex user-item interactions through multi-hop message passing. Nevertheless, these methods often fail to account for the dynamic nature of user-item interactions, leading to challenges in adapting to changes in user preferences and the distribution of new data. Consequently, their scalability and performance in real-world dynamic settings are constrained. In our study, we introduce GraphPro, a framework that merges dynamic graph pre-training with prompt learning in a parameter-efficient manner. This innovative blend enables GNNs to adeptly grasp both enduring user preferences and transient behavior changes, thereby providing precise and up-to-date recommendations. GraphPro tackles the issue of changing user preferences through the integration of a temporal prompt mechanism and a graph-structural prompt learning mechanism into the pre-trained GNN architecture. The temporal prompt mechanism imprints time-related information onto user-item interactions, equipping the model to inherently assimilate temporal dynamics, while the graph-structural prompt learning mechanism allows for the application of pre-trained insights to new behavior dynamics without continuous retraining. We also introduce a dynamic evaluation framework for recommendations to better reflect real-world situations and narrow the offline-online discrepancy. Our comprehensive experiments, including deployment in a large-scale industrial context, demonstrate the effortless plug-in scalability of GraphPro alongside various leading recommenders, underscoring the superiority of GraphPro in effectiveness, robustness, and efficiency. The implementation details and source code of our GraphPro are available in the repository at https://github.com/HKUDS/GraphPro.
copyright: acmcopyright
@@ 1. Introduction
Recommender systems are integral to numerous Web platforms, assisting users in navigating through the overwhelming amount of information by suggesting relevant items. In recent years, graph neural networks (GNNs) have emerged as powerful tools for modeling user-item interactions in recommendation tasks, enabling effective representation learning on graph-structured data. By treating users and items as nodes and their interactions as edges, GNNs can capture intricate multi-hop relationships between users and items, facilitating the generation of personalized recommendation results.
Previous studies in GNN-augmented recommendation have centered on crafting effective message passing strategies to delineate the collaborative relations between users and items (Berg et al., 2017; Ying et al., 2018; Wang et al., 2019). These efforts seek to harness GCN’s potential to unravel high-order connections within the user-item interaction graph. Follow-up investigations have delved into streamlining the message-passing approach (Chen et al., 2020a; He et al., 2020), diminishing GNN model intricacies (Mao et al., 2021; Shen et al., 2021), and advancing sampling method quality (Huang et al., 2021). Lately, the field has progressed with the integration of self-supervised learning (SSL) into GNN frameworks for recommendation (Ren et al., 2023a). These cutting-edge methodologies (Wu et al., 2021; Cai et al., 2023a; Yu et al., 2022) predominantly utilize the InfoNCE loss function (Oord et al., 2018) to align contrastive perspectives, thus enhancing the performance of the LightGCN architecture (He et al., 2020).
Although these methods have demonstrated impressive performance, they have primarily focused on static scenarios (Figure 1 upper), disregarding the dynamic nature of recommendation. In real-world scenarios (Figure 1 lower), the recommendation system operates in a dynamic setting where the model continuously learns from newly arriving data and predicts for the current time (Xiao et al., 2023; Yang et al., 2023a). However, existing methods are mainly designed for single-graph training and evaluation, leading to a degradation in recommendation dynamics and widening the offline-online gap (Ji et al., 2023). Furthermore, the arrival of new data may exhibit distribution shifts, further complicating the task of making accurate recommendations for graph-based recommendation models without incorporating useful contextual information for the newly arrived data. These challenges significantly limit the scalability of existing models and hinder their ability to adapt to evolving user preferences in a timely manner. This is critical for providing up-to-date and precise recommendations in dynamic environments.
@@
To address these challenges, we present a simple and effective framework called GraphPro, which combines parameter-efficient and dynamic Graph Pre-Training with Prompt Learning for recommendation. Our method begins with pre-training Graph Neural Networks (GNNs) on extensive historical interaction data, enabling them to capture long-term user preferences and item relevance. This pre-training phase assimilates knowledge from a substantial amount of historical interactions. Subsequently, during the fine-tuning phase on more recent target data, our model swiftly adapts to evolving user preferences and captures short-term behavior dynamics. This is achieved through a prompt learning schema, which facilitates effective knowledge transfer.
@@
To ensure the effective handling of evolving user preferences by the pre-trained GNN, our GraphPro framework seamlessly integrates a temporal prompt mechanism and a graph-structural prompt learning mechanism. This integration allows for the injection of time-aware context from new data, enabling the model to adapt to changing user preferences. Inspired by advancements in relative positional encoding techniques (Press et al., 2021; Su et al., 2021), we meticulously design a dedicated temporal prompt mechanism that aligns with the message aggregation layer of GNNs. Within this prompt mechanism, we encode time information on interaction edges as part of the normalization term for aggregation, all in a parameter-free manner. This innate capability allows the model to naturally incorporate temporal information without the need for additional fine-tuning. By incorporating temporal awareness into the pre-trained graph model, we empower the model to effectively capture vital signals that are highly relevant to the evolving user preferences.
@@
Moreover, our graph-structural prompt learning mechanism facilitates the seamless transfer of knowledge from the pre-trained model to downstream recent recommendation tasks. This framework eliminates the need for continuous incremental learning of the pre-trained model, enabling the transfer of pre-trained knowledge to any future time period to effectively adapt to behavior dynamics. In this mechanism, we include the newly generated interaction edges between the fine-tuning time and the pre-training time as prompt edges, providing the pre-trained model with essential contextual information for the fine-tuning process. Rather than undergoing extensive training, we perform a single non-training forward pass on the prompt edges, prompting the pre-trained model to adapt to the distribution shift of node representations and effectively adjust its predictions. It’s worth mentioning that our GraphPro is model-agnostic and parameter-efficient, making it easy to integrate into existing GNN recommenders as a plug-in enhancement. In summary, the main contributions of our work are as follows:
@@
  • We emphasize the importance of effectively and scalably pre-training and fine-tuning graph-based recommenders to accommodate time-evolving user preferences, thus facilitating up-to-date and accurate recommendations in dynamic environments.
@@
  • We introduce GraphPro, which effectively handles evolving user preferences through the pre-training and fine-tuning of GNNs. The proposed prompt learning paradigm enables the transfer of relevant knowledge from the pre-trained model to downstream recommendation tasks in both temporal and structural ways.
@@
  • Furthermore, we introduce a snapshot-based dynamic setting for recommender system evaluation, which offers a more realistic approximation to real-world recommendation scenarios compared to the traditional single-time testing approach.
@@
  • We conduct experiments on diverse datasets to showcase the robustness, efficiency, and performance advantages of GraphPro. To further validate the effectiveness of our framework, we also present an industry deployment on a large-scale online platform.
@@ 3.1.2. Temporal Prompt-enhanced Graph Convolutions
Here, $t^{\text{unix}}_{u,v}$ is the Unix timestamp assigned to the edge $e_{u,v}$, and the $\lfloor$*$\rfloor$ * ⌋ notation denotes the floor operation. To avoid the influence of specific numerical scales and ensure uniformity, we normalize these time attributes $\boldsymbol{t}={t_{u,v}|e_{u,v}=1}$ to the range of [0, 1].
@@
3.2. Fine-Tuning with Graph
@@ 4. Evaluation
In this section, we compare our proposed GraphPro with state-of-the-art methods to address the following research questions.
  • RQ1: Can GraphPro outperform other time-aware graph learning models and pre-trained GNNs in dynamic recommendations?
@@
  • RQ3: Can GraphPro perform equally well or even better than the vanilla full-data training paradigm?
@@ 4.1.1. Datasets
We utilize three public datasets that cover diverse real-world scenarios in dynamic recommendation. The Taobao dataset captures implicit feedback from Taobao.com, a Chinese e-commerce platform, over a period of 10 days. The Koubei dataset, provided for the IJCAI’16 contest, records 9 weeks of user interactions with nearby stores on Koubei in Alipay. The Amazon dataset consists of a 13-week collection of product reviews sourced from Amazon. More details about these datasets can be found in Table 1.
@@ 4.1.2. Baseline Models
We include the recent dynamic GNNs and graph prompt approaches as our baselines. Specifically, three most relevant research lines are included for comparison:
Dynamic Recommendation Methods. We include DGCN (Li et al., 2020), which explores categorizes edges as past and current and designing a GNN to propagate the information from past edges to current ones. However, it does not explicitly consider a dynamic setting with snapshots and focuses on a single graph.
@@
Dynamic Graph Neural Networks. These networks are tailored to dynamic graphs, updating embeddings with time sensitivity to reflect graph changes. We benchmark our approach against notable models EvolveGCN (Pareja et al., 2020) and ROLAND (You et al., 2022):
  • EvolveGCN (Pareja et al., 2020). It adapts to graph evolution using recurrent neural networks to modify GCN parameters over time, available as hidden state (-H) or recurrent input (-O) variants.
  • ROLAND (You et al., 2022). A dynamic graph method that employs meta-learning to refresh embeddings for subsequent re-initialization, integrating these with GNN’s layer-wise hidden states.
@@ 4.1.3. Integration with GNN Recommenders
GraphPro is a versatile architecture that seamlessly integrates as a plug-in component with any GNN-based recommender, highlighting its flexibility. In our evaluation, we implement GraphPro using the efficient LightGCN (He et al., 2020) model. Additionally, we extend GraphPro’s applicability by integrating it with self-supervised learning-based recommenders like SGL (Wu et al., 2021), MixGCF (Huang et al., 2021), and SimGCL (Yu et al., 2022). This integration provides empirical evidence of GraphPro’s effectiveness in enhancing dynamic and adaptable recommendations.
@@ 4.1.4. Evaluation Protocols
In our evaluation, we model real-world dynamics using graph snapshots at varying intervals, such as daily or weekly. A two-step sliding window technique is applied to learn from current data and predict subsequent changes. Under the Pre-train and Fine-tune framework, we pre-train with a substantial dataset fraction, fine-tune, and test on the latter snapshots (refer to Table 1). Baselines adopt the same approach to maintain consistency. Dynamic GNNs start fine-tuning with weights from the pre-training phase. We average our results over all future temporal snapshots, using standard metrics such as Recall@k and nDCG@k at k=20, in line with existing methodologies (He et al., 2020; Wu et al., 2021; He et al., 2017).
@@ 4.2.1. Comparison with Baselines
We showcase the performance of GraphPro alongside other methods, such as graph prompt techniques and dynamic GNNs, using LightGCN as the baseline, detailed in Table 2. Key observations from the analysis include:
  • Our GraphPro surpasses both graph prompt and dynamic graph learning methods, highlighting the efficacy of our pre-training and prompt learning strategy. This performance can be credited to: 1) Our temporal prompt mechanism that adeptly captures the evolving user-item interactions during pre-training and fine-tuning, and 2) Our graph prompt design that ensures seamless knowledge transfer from the pre-trained model and mitigates distribution shifts across temporal snapshots.
  • The varying performance of baseline methods underscores the complexity of dynamic recommendations. EvolveGCN excels with the Taobao dataset but not universally, possibly due to overfitting short-term user behaviors with its complex architecture. In contrast, GraphPro employs streamlined prompt mechanisms that adeptly seize long-term user interests and assimilate fresh behavioral data, offering a more consistent and effective approach.
@@ 4.2.2. Integration with SOTA Methods
We also evaluate GraphPro’s versatility with various base recommenders—MixGCF, SGL, and SimGCL—reimplementing all methods under uniform evaluation criteria. The averaged results across time slots can be found in Table 2. Our summarized observations are as follows:
  • GraphPro consistently excels alongside advanced recommenders, showcasing our method’s flexibility and capability to elevate performance across scenarios. Baselines, however, show varied performance depending on the base recommender and dataset used, with no single method dominating across all scenarios. While EvolveGCN-O often ranks second on Taobao, GPF stands out on Amazon regardless of the base recommender. Such variability in response to dataset specifics and underlying models prevents baselines from achieving consistent, substantial outcomes.
  • Superior base model representations typically yield better results. GraphPro sees a 4.5% boost when comparing SGL to SimGCL on Taobao, suggesting it capitalizes well on the base model’s improved representations in dynamic contexts. Conversely, methods like EvolveGCN-O fail to show gains with better base models, hinting at potential generalization weaknesses in their design.